# Discovery and Expansion of New Domains within Diffusion Models

### v2

@@ F.2.2 Discussion on Model Designs
Among four base DDPMs we have tested, there are two architecture variants namely the improved DDPM [36] and vanilla DDPM [19]. The difference between the two variants lies within the scheduler design for the Gaussian perturbation kernels: improved DDPM uses a cosine scheduler while vanilla DDPM adopts a linear one. Our experiments suggest that iDDPM in general synthesizes images with better quality in terms of FID scores, which aligns with previous studies [36, 74]. One implication from the above observation is that the domain generalization abilities studied in this context is inherited from the performance of model’s original performance.
F.2.3 More Qualitative Results
We show more qualitative samples from the CLIP-tuned methods in Fig. 7 and Fig. 8. More synthesized examples of our proposed method are included in Fig. 9. We also show part of the raw natural image samples used in our work in Fig. 10, Fig. 11, and Fig. 12, which helps to evaluate the diversity of the generated data.
Figure 7: Additional qualitative results from the fine-tuned DDPMs on different domains. We show qualitative examples for the domain human faces, bedrooms and churches generated from CLIP fine-tuned diffusion models using the same amount of OOD images (1K).
Figure 8: Additional qualitative results from the CLIP fine-tuned DDPMs on the astrophysical data.
Figure 9: Additional qualitative results from our proposed method for synthesizing OOD data without tuning the model parameters.
Figure 10: Examples of raw human face images used as OOD samples.
Figure 11: Examples of raw church images used as OOD samples.
Figure 12: Examples of raw bedroom images used as OOD samples.

### v3

@@ E.2.2 Discussion on Model Designs
Among four base DDPMs we have tested, there are two architecture variants namely the improved DDPM [nichol2021improveddmp] and vanilla DDPM [ho2020dpm]. The difference between the two variants lies within the scheduler design for the Gaussian perturbation kernels: improved DDPM uses a cosine scheduler while vanilla DDPM adopts a linear one. Our experiments suggest that iDDPM in general synthesizes images with better quality in terms of FID scores, which aligns with previous studies [nichol2021improveddmp, zhu2023boundary]. One implication from the above observation is that the domain generalization abilities studied in this context is inherited from the performance of model’s original performance.
E.2.3 More Qualitative Results
We show more qualitative samples from the CLIP-tuned methods in Fig. 6 and note that the tuning-based methods often fail to generalize to new image domains with large distributional shofts.
Figure 6: Fine-tuning methods often fail to transfer the original trained domain to the target unseen domain with large distributional shifts. We qualitatively show how a given ID sample (e.g., a dog RGB image) changes as the tuning epoch increases, using extra CLIP semantic guidance.
In Fig. 7, we first show the correlation between the diversity among generated samples with respect to the number of raw samples from the new target domains used in our proposed method. Specifically, we measure the diversity using the LPIPS scores, and note two takeaway information: First, the diversity increases when more samples are available, and our empirical findings suggest that N=8001000N=800-1000 is a reasonable choice. Next, the generated data from our proposed sampling method exhibits reasonable diversity among samples within the same target domain. Finally, it is important to note that the LPIPS score does not directly reflect the quality of the generated images. For example, in the case of galaxy images under large domain shift, the LPIPS score may appear lower despite the good perceptual quality of the outputs, since such images often share similar dark backgrounds and structural patterns by nature.
Figure 7: Effect of the number of new samples from the target domains on generation diversity. Results are shown for two OOD domains: Cat (small distribution gap) and Galaxy (large distribution gap) via diffusion model pre-trained on dogs.
We also qualitatively visualize the generated data between the interpolation between two raw samples in Fig. 8. Overall, the results demonstrate noticeable visual variation across the full interpolation path, with higher similarity observed between samples that are spatially closer in the latent space. In practice, we typically select the samples from the middle of the interpolation, as it offers a better trade-off between quality and diversity.
Figure 8: Visualization of the interpolation between two raw samples. Uncurated samples are shown for two OOD domains: Cat (small distribution gap) and Galaxy (large distribution gap) via diffusion model pre-trained on dogs.
Figure 9: More uncurated samples for various OOD domains generated via our proposed method from frozen diffusion models.
Finally, Fig. 9 includes more uncurated generation results from our proposed sampling-centric method from frozen pre-trained unconditional diffusion models.